QUESTION 25

Suppose that X is a binomial random variable with n D 25 and p D 0:5: Calculate

(a) the probability P .X D 15/: (2)

(b) the probability P .X  16/: (4)

(c) the expected value of X. (2)

(d) the variance of X. (3)



1
Expert's answer
2022-05-20T11:44:50-0400

"X\sim Bin (25, 0.5)"

(a)


"P(X=15)=\dbinom{25}{15}(0.5)^{15}(1-0.5)^{25-15}"

"=0.09741663933"

(b)


"P(X>16)=P(X=17)+P(X=18)"

"+P(X=19)+P(X=20)"

"+P(X=21)+P(X=22)"

"+P(X=23)+P(X=24)"

"+P(X=25)"

"=\dbinom{25}{17}(0.5)^{17}(1-0.5)^{25-17}"

"+\dbinom{25}{18}(0.5)^{18}(1-0.5)^{25-18}"

"+\dbinom{25}{19}(0.5)^{19}(1-0.5)^{25-19}"

"+\dbinom{25}{20}(0.5)^{25}(1-0.5)^{25-20}"

"+\dbinom{25}{21}(0.5)^{21}(1-0.5)^{25-21}"

"+\dbinom{25}{22}(0.5)^{22}(1-0.5)^{25-22}"

"+\dbinom{25}{23}(0.5)^{23}(1-0.5)^{25-23}"

"+\dbinom{25}{24}(0.5)^{23}(1-0.5)^{25-24}"

"+\dbinom{25}{25}(0.5)^{23}(1-0.5)^{25-25}"

"=0.05387607217"

"P(X\ge16)=P(X=16)+P(X>16)"

"=\dbinom{25}{16}(0.5)^{16}(1-0.5)^{25-16}+P(X>16)"

"=0.11476147175"

(c)


"E(X)=np=25(0.5)=12.5"



(d)


"Var(X)=np(1-p)=25(0.5)(1-0.5)=6.25"